### v1

# A detailed study on ergodic to non-ergodic phase transition in the dissipative anisotropic Dicke model

Abstract
Ergodic-non-ergodic phase transition is one of the paramount features of the anisotropic Dicke model. Here, we have thoroughly examined the effect of the dissipation by analyzing both the eigenvalue and eigenvector properties of the Liouvillian with the aid of the scaling of Liouvilian gap, and the average participation ratio. We show that the properties of the eigenvectors of Liouvillian are consistent with those of the eigenvalues, revealing a phase diagram, which has similarities to the non-ergodic to ergodic transition in the closed undriven system. We also uncover that the Liouvillian gap is independent of system size in the non-ergodic phase whereas in the ergodic phase, it scales with the atom number as: $N^{-z}$ where $10<z<$ 1. Moreover, we extend our analysis to the driven case where a Thue-Morse quasi-periodic sequence is applied and observe that the boson dissipation plays a pivotal role in stabilizing the prethermal plateau. Our investigation indicates that a non-ergodic phase is more favorable than the ergodic phase in the presence of bosonic dissipation.
I Introduction
Exploring the many-body quantum dynamics through atom-photon interaction is the heart of the atomic, condensed matter, and ultra-cold quantum physics Cohen-Tannoudji et al. (1998); Gutzler et al. (2021); Chang et al. (2018). Experiments using cavity QED setup provide an ideal platform to investigate the underlying microscopic properties e.g., the collective effects, various phase transitions, Bose-Einstein condensation, etc Walther et al. (2006); Reiserer and Rempe (2015); Vaidya et al. (2018). From a theoretical perspective, the Dicke model (DM) is an iconic model for understanding the above aspects of such systems Garraway (2011); Kirton et al. (2019); Villaseñor et al. (2024).
The Dicke model describes the interaction between $N$ ensemble of atoms with a single mode bosonic field via a dipolar coupling strength  Dicke (1954); Lambert et al. (2004); Emary and Brandes (2003a, b). In the thermodynamic limit ($N\to\infty$) the model shows a quantum phase transition from normal to superradiant phase Dicke (1954); Chávez-Carlos et al. (2016); Kirton and Keeling (2018); Kirton et al. (2019); Das and Sharma (2022). Along with it, such a model also shows the excited state phase transition(ESQPT), and thermal phase transition(TPT) under different circumstances. A generalized version of the Dicke model namely the anisotropic Dicke model (ADM) Buijsman et al. (2017); Kloc et al. (2017); Shapiro et al. (2020); Das et al. (2023a), where the coupling strengths correspond to the rotating and counter-rotating terms are considered to be different. Due to the asymmetry in the coupling strengths, ADM shows novel features in addition to the well-known existing properties, i.e., ergodic to non-ergodic transition (ENET) Buijsman et al. (2017); Hu and Wan (2021). There exists a one-to-one correspondence of the non-ergodic phase to the integrability of the system. Hence, a detailed study on the non-ergodic phase will be useful in understanding the following emerging phenomena, e.g., many-body localization, quantum scars, prethermalization, etc Polkovnikov et al. (2011); Mori et al. (2018); Serbyn et al. (2021). On the other hand, the ergodic phase connects to the non-integrable phase, a path to study the quantum chaos D’Alessio et al. (2016); Borgonovi et al. (2016). Generally, the question arises regarding the behavior of the above-mentioned phases in the presence of dissipation.
In most cases, a leaky cavity is considered to introduce the dissipative effects on the atom-photon interaction Bhaseen et al. (2012); Kirton et al. (2019); Kirton and Keeling (2017, 2018). It was shown that such an effect is important for understanding the self-organization and discrete time-crystalline order in this model Keeling et al. (2010); Gong et al. (2018). In the case of superradiant phase transition, the effect of dissipation is previously explored Kirton and Keeling (2018). It is noteworthy to study the impact of dissipation in the ENET, which we explore in this manuscript. The Hamiltonian approach is insufficient to capture such dissipative dynamics. Hence, we adapt the Liouvillian approach motivated by the Lindblad quantum master equation to incorporate the non-unitary dynamics Breuer et al. (2002).
In this prescription, the Liouvillian gap ($\Delta$) plays an important role in characterizing the asymptotic convergence to the steady state Haga et al. (2021); Zhou et al. (2022); Prosen and Pižorn (2008); Kessler et al. (2012); Cai and Barthel (2013); Kastoryano and Eisert (2013); Žnidarič (2015); Casteels et al. (2017); Minganti et al. (2018); Shibata and Katsura (2019a, b, 2020); Mori and Shirai (2020, 2023); Shirai and Mori (2023); Mori (2024); Landi et al. (2022); Yuan et al. (2021), and one of our main focuses here is to study its dependence on the system size in different phases of this model. Recently, it was shown that for the boundary-dissipated one-dimensional quasiperiodic and disordered systems Zhou et al. (2022), $\Delta$ scales as exponentially for localized phase and $L^{-3}$ (where $L$ is the system size) for extended phase. For other types of boundary-dissipated systems, numerical results have shown that $\Delta$ scales with $L^{-z}$, where $z\in[1,2]$ for chaotic systems and $z=3$ for integrable systems  Prosen and Pižorn (2008); Žnidarič (2015); Shibata and Katsura (2020); Mori and Shirai (2020). On the other hand, for a special case ( XXZ chain in the presence of bulk dissipation ), it was found that, $\Delta$ is independent of the system size ($L$) for the ferromagnetic case, whereas it decays as $L^{-0.8}$ in the antiferromagnetic case Žnidarič (2015). Here, we study the behavior of $\Delta$ as a function of $L$ for dissipative ADM to characterize different phases. Moreover, we also investigate the properties of eigenvectors of the Liouvillian and compare their consistency with the corresponding eigenvalues.
Another important aspect of a many-body quantum system is exploring the dynamics under an external drive Eckardt (2017). In such cases, it is expected that the system will reach the infinite temperature state (e.g., thermal state) Abanin et al. (2019); Nandkishore and Huse (2015). An important distinction can be observed for the periodic drive, where instead of thermalization, a prethermal quasi-steady state emerges in the intermediate time scale Abanin et al. (2015); Beatrez et al. (2021); Saha and Bhattacharyya (2023); Das et al. (2023b). Using the well-celebrated Floquet theory, an effective Hamiltonian can be constructed for the prethermal state in high-frequency limit and the relevant conserved quantities can be derived subsequently Fleckenstein and Bukov (2021). More recently, research interest shifted towards exploring the quantum dynamics under quasi-periodic drive Nandy et al. (2017); Mukherjee et al. (2020); Zhao et al. (2021, 2022); Das et al. (2023b); Tiwari et al. (2024, 2025), where it is expected that the lifetime of the prethermal phase is shortened due to the absence of temporal order. As such, the quasi-periodic drive destroys the integrability of the system and pushes it to the thermal state. A similar study was also reported in our previous work of ADM Das et al. (2023b). In this manuscript, we extend this analysis to the dissipative case, where we study the interplay between the quasi-periodic drive and dissipation and its effect on the stability of the prethermal state.
The structure and the key points of the manuscript are discussed below. A brief description of the dissipative ADM (i.e., a spin-boson system interacts with another bosonic environment) is given in section-II. The dynamical equation of the system is given in Eq. (3). We also discuss the criticality in the system and the effect of dissipation on the critical line in section- II.1. In section- III, we study the scaling of $\Delta$ for our system, as a function of the total Liouvillian space dimension ($N_{\text{L}}$) and the atom number ($N$), separately for the nonergodic and the ergodic phase. Furthermore, we study the eigenvector property with the aid of the average participation ratio for left, right, and both left-right eigenvectors separately in the same section.
The effect of an external drive on dissipative ADM is provided in section IV, where we consider a quasiperiodic drive modeled by the Thue-Morse sequence Thue (1906); Nandy et al. (2017); Mukherjee et al. (2020); Zhao et al. (2021, 2022) and study the quantum dynamics with the aid of average boson number and the mutual information between spins and bosons in the system. We mostly focus on the changes in the behavior of our model in comparison to its corresponding closed version, which has not been explored earlier. Finally, we conclude in section-V and discuss the broader perspective of our results.
Figure 1: Schematic picture for the dissipative Dicke model, where the system is defined by the interaction between an ensemble of $N$ two level atoms with a single mode ($\omega$) leaky bosonic cavity and the whole system is connected with another bosonic environment. The gap between the two levels of each atom is $\omega_{0}$ and the bosonic dissipation strength is $\kappa$.
II Description of the system
Here we consider a system consisting of $N$ non-interacting two-level atoms coupled with a single-mode bosonic leaky cavity. The corresponding Hamiltonian of the system is:
$\mathcal{H}$ $=$ $\omega a^{\dagger}a+\omega_{0}J_{z}+\frac{g_{1}}{\sqrt{N}}(a^{\dagger}J_{-}+aJ_{+})+$ (1)
$\frac{g_{2}}{\sqrt{N}}(a^{\dagger}J_{+}+aJ_{-}).$
where $g_{1}$ and $g_{2}$ are the rotating and counter-rotating spin-boson coupling strengths as we are interested in the anisotropic Dicke model Das et al. (2023a). Here $a$ ($a^{\dagger}$) is the bosonic annihilation (creation) operator, and $J_{\alpha}=\sum_{i=1}^{N}\frac{1}{2}\sigma_{\alpha}^{i}$ ($\alpha=$ $x$, $y$ $z$), total angular momentum of the ensemble of $N$ spins. $\omega$ is the bosonic mode frequency and $\omega_{0}$ is the level splitting of the atoms. The first two terms represent the free Hamiltonian of the bosonic cavity and atoms respectively. The basis of the Hamiltonian is: $|n,m\rangle$, where $n=0,1,2...,n_{\text{max}}$: boson number and for symmetric spin subspace considering even $J$, $m=-J,-J+1,...,0,...,J-1,J$. Thus the Hilbert space dimension of the closed system is: $N_{\text{D}}=(N+1)(n_{\text{max}}+1)$. $g_{1}=g_{2}=g$ recovers the original isotropic Dicke model.
We assume that the decay rate of the leaky cavity is $\kappa$ and all the other sources of dissipation and dephasing are excluded in our model. The coupling between the cavity and the environment is weak and it can be modeled using Jaynes-Cummings Hamiltonian Breuer et al. (2002). The presence of the environment leads to dissipative dynamics which can be explained by Lindblad master equation in the Markovian regime Breuer et al. (2002). In Fig. 1 we show a schematic diagram of the dissipative isotropic Dicke model.
The dynamical equation of atom and leaky cavity is given by:
$\frac{d\rho}{dt}=\mathcal{L}\rho,$ (2)
where $\rho$ is total density matrix of the total system and $\mathcal{L}$ is the Liouvillian super-operator. The explicit form of $\mathcal{L}\rho$ is:
$\mathcal{L}\rho=-i[H,\rho]+\kappa[2a\rho a^{\dagger}-\{a^{\dagger}a,\rho\}].$ (3)
The corresponding analytical form of $\mathcal{L}$ is given by:
$\mathcal{L}$ $=$ $-i\left[(\mathcal{H}\otimes\mathcal{I})-(\mathcal{I}\otimes\mathcal{H}^{T})\right]+\kappa\Big{(}2a\otimes a^{*}-a^{\dagger}a\otimes\mathcal{I}$ (4)
$-\mathcal{I}\otimes a^{\dagger}a\Big{)}.$
The basis of the Liouvillian is $|n_{l},m_{l}\rangle\otimes|n_{r},m_{r}\rangle$ and hence the Liouvillian space dimension is $N_{\text{L}}=N_{\text{D}}\times N_{\text{D}}$. Using parity symmetry, the dimension of the Liouvillian space reduces to $\frac{N_{\text{ L}}}{2}$ Prasad et al. (2022).
II.1 Existence of the critical points
Criticality in the closed Dicke model is well known Das and Sharma (2022) and in the presence of anisotropy, how this critical point is generalized to a critical line is also explored previously Das et al. (2023a). In this manuscript, one of our main goals is to understand the effect of dissipation on the critical line. For the bosonic dissipation, the relation between the critical coupling parameters for the second-order phase transition is given by Kirton et al. (2019):
$\left(g_{1}^{2}-g_{2}^{2}\right)^{2}-2\left(g_{1}^{2}+g_{2}^{2}\right)\omega\omega_{0}+\left(\omega^{2}+\kappa^{2}\right)\omega_{0}^{2}=0,$ (5)
which separates the non-ergodic and ergodic phases. In case of isotropic Dicke model ($g_{1}=g_{2}=g$), the equation will reduce to: $g^{*}=\frac{1}{2}\sqrt{\omega\omega_{0}}\sqrt{1+\kappa^{2}/\omega^{2}}$. For the numerical simplicity, we choose $\omega=\omega_{0}=1$. To include the dissipation we choose $\kappa=1$ and hence the critical point for the quantum phase transition from integrable to chaotic phase is: $g^{*}=\frac{1}{\sqrt{2}}\approx 0.7071$.
If we gradually increase the value of $\kappa$, then the regime of the non-ergodic phase will also be extended which means the non-ergodic phase is more favorable under the bosonic dissipation. In the next section, we quantify the above result by studying several properties of the Liouvillian.
III Results
III.1 Finite size scaling of the Liouvillian gap
Similar to the closed model, dissipative ADM also shows a transition from non-ergodic to ergodic phases (see Fig. 9 in the Appendix). In this section, we study the eigenvalue property of the Liouvillian $\mathcal{L}$, the scaling of the Liouvillian gap (as a function of the total Liouvillian space dimension) in different regions of the phase diagram covering the ergodic and non-ergodic phases.
Figure 2: The Liouvillian gap as a function of (a) the atom number $N$ and (b) the total dimension $N_{\text{L}}$ of the Liouvillian space for $g_{1}=1.25$, and $g_{2}=0.1-1.5$ such that we cover both the nonergodic and the ergodic phase of the system. For each pair of $(g_{1},\ g_{2})$ the range of the atom number is $N=4-10$. In the nonergodic phase $\Delta$ is independent of the system size whereas, in the ergodic phase: (a) $\Delta\propto N^{z}$. For, $g_{2}=1.0:\ \ln\Delta=-0.42[\ln N]-0.0435$, $g_{2}=1.1:\ \ln\Delta=-0.578[\ln N]+0.124$, $g_{2}=1.2:\ \ln\Delta=-0.75[\ln N]+0.28$, $g_{2}=1.25:\ \ln\Delta=-0.84[\ln N]+0.362$, $g_{2}=1.3:\ \ln\Delta=-0.897[\ln N]+0.38$, $g_{2}=1.4:\ \ln\Delta=-0.9945[\ln N]+0.384$, and $g_{2}=1.5:\ \ln\Delta=-1.053[\ln N]+0.34$; (b) $\Delta\propto N_{L}^{z}$ (in the ergodic phase). For $g_{2}=1.0:\ \ln\Delta=-0.244[\ln N_{L}]+1.76$, $g_{2}=1.1:\ \ln\Delta=-0.336[\ln N_{L}]+2.609$, $g_{2}=1.2:\ \ln\Delta=-0.445[\ln N_{L}]+3.61$, $g_{2}=1.25:\ \ln\Delta=-0.514[\ln N_{L}]+4.27$, $g_{2}=1.3:\ \ln\Delta=-0.55[\ln N_{L}]+4.56$, $g_{2}=1.4:\ \ln\Delta=-0.61[\ln N_{L}]+5.023$, and $g_{2}=1.5:\ \ln\Delta=-0.65[\in N_{L}]+5.26$. For all the plots $\omega=\omega_{0}=1$, $\kappa=1$. The bosonic cut-off: $n_{\text{max}}=26$.
Figure 3: Average of participation ratio ($P_{\text{avg}}$), using (a) the left (b) right, and (c) biorthonormal basis considering both left and right eigenvectors of the Liouvillian for dissipative ADM as a function of the rotating and counter-rotating coupling parameters $g_{1}$ and $g_{2}$. The atom number: $N=10$, the bosonic cut-off: $n_{\text{max}}=26$. For all the figures $\omega_{0}=\omega=1,\ \kappa=1$.
Figure 4: Average of participation ratio ($P_{\text{avg}}$), using both the left and right eigenvectors of the Liouvillian for dissipative DM as a function of the coupling parameter $g$. The atom number: $N=10$, the bosonic cut-off: $n_{\text{max}}=26$. $\omega=\omega_{0}=1$, $\kappa=1$.
The Liouvillian gap, $\Delta$ is defined by the real distance of the nearest neighbor eigenvalue of $\mathcal{L}$ from its zero eigenvalue (eigenvalue corresponds to the steady state):
$\Delta=\text{min}[\text{Re}(E_{0}-E_{n})],$ (6)
where $E_{n}$ are the eigenvalues of the liouvillian and $E_{0}=0+0i$, is the corresponding minimum eigenvalue. In Fig. 2 we study the Liouvillian gap as a function of the total dimension of the Liouvillian space for $g_{1}=1.25$, and $g_{2}=0.1-1.5$, ranging from the non-ergodic/regular phase to the ergodic/quantum chaotic phase. We notice that in the ergodic phase, $\Delta\propto N^{-z}_{L}$ with $10<z<$ 1, whereas in the non-ergodic phase it is independent of the total Liouvillian space dimension. We also plot the same data for $\Delta$ as a function of the atom number $N$, (for the fixed $n_{\text{max}}$) and see a similar behavior (scaling) as the previous case. We note that our scaling result is similar to one of the cases of bulk dissipation highlighted recently in the work by Znidaric. Žnidarič (2015). The reason for this type of scaling is given below.
In the ergodic phase, the system dynamics occurs in the full Hilbert space, so by increasing the number of atoms, the time to cover the full space also increases. The negative slope in the plots indicates that the relaxation time increases with the atom number. On the other hand, in the case of the non-ergodic phase, due to the presence of symmetry and conserved quantities, the dynamics takes part in a relatively smaller subspace. Hence for a higher number of atoms, no such changes in the $\Delta$ is noticed here. We consider only the spontaneous emission process in the dissipative ADM and hence it shows different behavior from the result of the previous work considering the dephasing Mori (2024).
However, we confine ourselves to the finite number of atoms and observe that in this limit, the system has a finite $\Delta$. Our future focus will be on the behavior of ($\Delta$) in the thermodynamic limit.
III.2 Eigenstate properties
The eigenvalue properties of the dissipative Dicke model are well studied Prasad et al. (2022). We also provide a brief discussion of the eigenvalue properties in the Appendix. A recent study exists in the literature that includes the anisotropic case of this particular model Villaseñor et al. (2024). However, a complete study requires both the eigenvalue and eigenvector properties to understand the phase diagram of this model. In this section, we discuss the eigenstate property and check its consistency with the eigenvalue property.
III.2.1 Average participation ratio
It is well known that analyzing the eigenstates of non-Hermitian systems can predict the emergence of chaos, e.g., many-body coupled Hatano-Nelson chain Suthar et al. (2022). Motivated by the above approach, we study the eigenstate properties of the Lindbladian super-operator. In particular, we focus on the average participation ratio (PR), $P_{\text{avg}}$. Similar to the non-Hermitian Hamiltonian, the right and left eigenstates are not the same. Therefore, we perform a bi-orthogonalization to make an orthogonal basis that can span the entire Liouvillian space, which is given below in detail. We also note that the eigenstates with negative eigenvalues have a finite lifetime Saha and Bhattacharyya (2024). Hence, our results can only be envisaged as the chaotic properties of the non-equilibrium states which may not have an equilibrium counterpart.
In the case of a Liouville space, we perform the LU decomposition for bi-orthogonalization. Let $L$ and $R$ be the corresponding left and right eigenvectors of $\mathcal{L}$ and $L^{\dagger}R=M$. The matrix $M$ can be represented in its LU-decomposed form: $M=M_{\text{L}}M_{\text{U}}$, where $M_{\text{L}}(M_{\text{U}})$ is the lower(upper) triangular matrix. Hence,
$L^{\dagger}R=M_{\text{L}}M_{\text{U}}$
or
$M_{\text{L}}^{-1}L^{\dagger}RM_{\text{U}}^{-1}=1.$ (7)
Following the above approach, one can redefine the left and right eigenvectors as $L\rightarrow L\left(M_{\text{L}}^{-1}\right)^{\dagger}$ and $R\rightarrow RM_{\text{U}}^{-1}$ such that $L^{\dagger}R=1$ to make the Liouvillian space bi-orthogonal.
For non-Hermitian systems, PR can be defined in two ways: first, using the $n$th left or right eigenstate, and the second is defined using the biorthogonal basis from both left and right eigenstates. The definition of PR is given by:
$P_{n}=\frac{\left(\sum_{i}|\psi_{i}^{n}|\right)^{2}}{\sum_{i}|\psi_{i}^{n}|^{2}}$ (8)
$P_{nB}=\frac{\left(\sum_{i}|\tilde{\psi}_{i}^{n}|\right)^{2}}{\sum_{i}|\tilde{\psi}_{i}^{n}|^{2}}.$ (9)
The subscript $nB$ represents biorthogonal PR, $\psi_{i}^{n}=(\langle R_{n}|b_{i}\rangle^{*})\langle R_{n}|b_{i}\rangle$ and $\tilde{\psi}_{i}^{n}=(\langle L_{n}|b_{i}\rangle^{*})\langle R_{n}|b_{i}\rangle$ with $R_{n}$ and $L_{n}$ are the $n^{\text{th}}$ right and left eigenstates respectively, and $|b_{i}\rangle$ are Fock space chosen as computational basis. The average PR is defined as: $P_{\text{avg}}=\sum_{n}P_{\text{n}}/N_{\text{L}}$ and $P^{B}_{\text{avg}}=\sum_{n}P_{\text{nB}}/N_{\text{L}}$. In Fig. 3 we show the value of $P_{\text{avg}}$ and $P_{\text{avg}}^{B}$ as a function of the coupling strengths. In the dark portion, the average participation ratio is relatively low which indicates the non-ergodic phase, compared to the light color regime where the average participation is quite high and represents the ergodic phase. The phase diagrams are similar to Fig. 9 in the Appendix, where we show the transition from the non-ergodic to the ergodic phase with the aid of level spacing ratio. Hence we can say that the eigenvector properties of $\mathcal{L}$ in our system are consistent with the eigenvalue properties. In Fig. 4 we highlight the data along the diagonal direction $g_{1}=g_{2}=g$ which indicates the isotropic Dicke model. Here the vertical dot-dashed line represents the critical coupling strength of the phase transition at $g=g^{*}=0.7071$. The variation between the three different types of participation ratios in Fig. 3 is interesting and worthy of a closer investigation.
Figure 5: (a, b, e, f, i, j, m, n) Average boson number, $N_{\text{av}}(t)$, and (c, d, g, h, k, l, o, p) mutual information between spins and bosons, $I_{\text{spin-boson}}$, as a function of time $t_{n}=2^{n}T$ for the dissipative ADM under the Thue-Morse quasiperiodic drive. The initial states have low energies, so that $\langle E_{\text{in}}\rangle=0.25$. Panels (a, c, e, g, i, k, m, o) represent the data for a fixed driving frequency $\omega_{\text{d}}=50$, $g_{1}=1.25$, and various values of $g_{2}$. Panels (b, d, f, h, j, l, n, p): $g_{1}=1.25$, $g_{2}=0.7$ and various values $\omega_{\text{d}}$. In all panels, the driving amplitude is $\Omega=1.0$, $N=6,\ n_{\text{max}}=20,\ \omega_{0}=\omega=1$. For (a-d) $\kappa=0.0$, (e-h) $\kappa=0.05$, (i-l) $\kappa=0.6$, (m-p) $\kappa=1.0$. The time t is in units of $\omega_{0}^{-1}$.
IV Analysis of the driven dissipative ADM
In our earlier study of driven closed ADM Das et al. (2023b), we show that under the external periodic drive, the system reaches a plateau that escapes Floquet heating. In contrast, for the quasiperiodic drive, the system features a prethermal plateau and finally reaches an infinite temperature state. Such analysis can also be extended for the dissipative case. From the analysis of the previous sections, we note that the non-ergodic phase becomes more stable in the presence of bosonic dissipation. Whereas, our previous evidence shows that the quasiperiodic drive pushes the system to the ergodic phase. Hence, initiating the competition between the quasiperiodic drive and the bosonic dissipation to stabilize the non-ergodic phase is the main focus of this section.
$Driving\ protocol\ -$ We briefly discuss the quasi-periodic driving protocol. In particular, we consider the quasi-periodic protocol consisting of the Thue-Morse sequence. Here, we define two Hamiltonian $H_{\pm}=\omega a^{\dagger}a+\omega_{0}J_{z}+\frac{(g_{1}\pm\Omega)}{\sqrt{N}}(aJ_{+}+a^{\dagger}J_{-})+\frac{(g_{2}\pm\Omega)}{\sqrt{N}}(aJ_{-}+a^{\dagger}J_{+})$, where $\Omega$ is the driving amplitude and the corresponding Liouvillian using Eq. 4 are $\mathcal{L}_{\pm}$. The Thue-Morse sequence Thue (1906); Nandy et al. (2017); Mukherjee et al. (2020); Zhao et al. (2021, 2022) can be constructed by starting time evolution operators $\mathcal{U}_{\pm}=\exp(\mathcal{L}_{\pm}T)$. We can build up the next sequence using $\mathcal{U}_{1}=\mathcal{U}_{-}\mathcal{U}_{+}$ and $\tilde{\mathcal{U}}_{1}=\mathcal{U}_{+}\mathcal{U}_{-}$, $\mathcal{U}_{2}=\tilde{\mathcal{U}}_{1}\mathcal{U}_{1}$ and $\tilde{\mathcal{U}}_{2}=\mathcal{U}_{1}\tilde{\mathcal{U}}_{1}$ and so on. The driving unit cells of time length $2^{n}T$ can be recursively constructed as: $\mathcal{U}_{n+1}=\tilde{\mathcal{U}}_{n}\mathcal{U}_{n},\ \tilde{\mathcal{U}}_{n+1}=\mathcal{U}_{n}\tilde{\mathcal{U}}_{n}$. The time evolution of the initial density matrix is given by $\rho_{n}|_{N_{\text{L}}\times 1}=\rho(t=2^{n}T)|_{N_{\text{L}}\times 1}=\mathcal{U}_{n}|_{N_{\text{L}}\times N_{\text{L}}}\times\rho_{\text{in}}|_{N_{\text{L}}\times 1}$, where $\rho_{\text{in}}|_{N_{\text{L}}\times 1}=|\psi_{\text{in}}\rangle|_{N_{\text{D}}\times 1}\otimes|\psi_{\text{in}}\rangle|_{N_{\text{D}}\times 1}$ is the initial density matrix and $|\psi_{\text{in}}\rangle$ is the initial state. To calculate the expectation value of an observable we reshape the density matrix as: $\rho(t)|_{N_{\text{L}}\times 1}\rightarrow\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}$. Here we take the average over $20$ low energy eigenstates (starting from the ground state) of the decoupled Hamiltonian ($g_{1}=g_{2}=0.0$).
IV.1 Quantum dynamics
To study quantum dynamics we consider the average boson number:
$N_{\text{av}}(t)=\text{Tr}\left[\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}\times(a^{\dagger}a)|_{N_{\text{D}}\times N_{\text{D}}}\right]$ (10)
and the mutual information between spins and bosons:
$I_{\text{spin-boson}}=S_{\text{spin}}+S_{\text{boson}}-S_{\text{spin-boson}},$ (11)
where $S_{\text{spin-boson}}=-\text{Tr}\left[\rho(t)\ln(\rho(t))\right]$, $S_{\text{spin(boson)}}=-\text{Tr}\left[\rho_{\text{spins(boson)}}(t)\ln(\rho_{\text{spins(boson)}}(t))\right]$ are the corresponding entropies and $\rho_{\text{spin}}(t)|_{(N+1)\times(N+1)}=\text{Tr}_{\text{boson}}\left[\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}\right]$, $\rho_{\text{boson}}(t)|_{(n_{\text{max}}+1)\times(n_{\text{max}}+1)}=\text{Tr}_{\text{spin}}\left[\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}\right]$ are the corresponding density matrix or reduced density matrix of the spin(boson) obtained by tracing over the bosonic(spin) degrees of freedom.
In Fig. 5, we consider low-energy initial states and the Thue-Morse driving protocol. We show the dynamics of the average boson number [Fig. 5(a, b, e, f, i, j, m, n)] and the mutual information between spins and bosons [Fig. 5(c, d, g, h, k, l, o, p)] for a fixed intermediate value of the driving frequency $\omega_{\text{d}}$ and various values of the coupling parameter $g_{2}$ [Fig. 5(a, c, e, g, i, k, m, o)]. We also study the same quantities for a fixed $(g_{1},g_{2})$ pair associated with the chaotic undriven model and various values of $\omega_{\text{d}}$ [Fig. 5(b, d, f, h, j, l, n, p)]. For the close system ($\kappa=0$) von Neumann entanglement entropy is a good quantifier as the spin-boson density matrix, $\rho(t)$ forms a pure state. On the other hand, when we switch on $\kappa$, $\rho(t)$ doesn’t remain a pure state, it becomes a mixed state. As mutual information is a good quantifier for analyzing a mixed state, we pick this observable for our study. Fig. 5(a-d) correspond to $\kappa=0$ and it shows a result similar to the closed quasiperiodically driven system. There are three characteristic features: an initial increasing part followed by an intermediate prethermal plateau and a finial infinite temperature state. On the other hand, all other panels with $\kappa\neq 0$ exhibit an initial increasing part, and then the system remains in a prethermal plateau which is not followed by the final infinite-temperature state, and this contrasts with the results in the closed ADM ($\kappa=0$ case), (see the Fig. 5(a-d) here, or the Fig. 4 in our earlier work Das et al. (2023b)). The reason behind the nature of the $\kappa\neq 0$ case is as follows: in a closed-driven system, the system absorbs energy from the drive and tends to reach an infinite-temperature state over time. At very high driving frequencies, this can lead to the emergence of a prethermal phase Das et al. (2023b). On the other hand, the bosonic dissipation leads the boson to the ground states, which lowers the entropy of the system.
For a better understanding of the effect of dissipation, we also provide the dynamical equation of the spins at the strong dissipation limit where the bosonic degrees of freedom can be adiabatically eliminated. For $\kappa>>g_{1},\ g_{2}$, the dynamical equation of the spins is given by Kessler et al. (2012); Link and Strunz (2020). ,
$\dot{\rho_{s}}=-i[\omega J_{\text{z}},\rho_{s}]+\kappa_{\rm eff}[2J_{-}\rho J_{+}-\{J_{+}J_{-},\rho_{s}\}].$ (12)
Here $\rho_{s}$ is the spin density matrix. The above dynamical equation represents the Dicke model of resonance fluorescence. We note that the quasi-periodic drive has no effect (in $\kappa>>g_{1},\ g_{2}$ limit) and the steady state will be the ground state.
Another interesting feature in the high $\kappa$ regime is the linear dependency of $\Delta$ on atom number $N$, which is well known for the Dicke superradiance burst and shows a sharp contrast from the finite values of $\kappa$ Gross and Haroche (1982).
V Conclusion
[…]
We also extend the analysis to the driven case where a quasiperiodic Thue-Morse sequence is applied to the system and study the quantum dynamics with the aid of average boson number and mutual information between spins and bosons in the system. Here we see the observables are initially increasing with the sequential time followed by a stabilized prethermal plateau. The system is not going to the final infinite temperature state which contrasts the result for the closed anisotropic quasiperiodically driven system and not reported previously.
Effect of the dissipation and dephasing in the superradiance phase transition was already studied in previous works by Kirton et al Kirton and Keeling (2017). Our work provides an addition to the field of the dissipative Dicke model, where the non-ergodic phase transition is thoroughly studied and analyzed for both the driven and undriven cases.
Further investigation considering the individual decay processes of the atoms is required to understand the other illuminating aspects of the dissipative Dicke model. The dimensionality of the Liouvillian is an issue for simulating large numbers of atoms, hence future research works demand advanced numerical techniques to overcome this difficulty Weimer et al. (2021).
Acknowledgments
We sincerely thank Auditya Sharma and Devendra Singh Bhakuni for fruitful discussions and constructive comments. We are also grateful to the High-Performance Computing (HPC) facility at IISER Bhopal, where large-scale calculations in this project were run. P.D. acknowledges IISERB for the PhD fellowship. This research was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the program Stability of Quantum Matter in and out of Equilibrium at Various Scales (ICTS/SQMVS2024/01).
Appendix A Eigenvalue properties
We give a review of the distribution function of the eigenvalues and levels spacings ratio of a many-body quantum system, which motivates us to study the nature of such function for this model. Based on the eigenvalue distribution of a Hermitian system, it can be broadly classified into two categories: (i) regular/integrable/non-ergodic and (ii) chaotic/ergodic. Berry and Tabor hypothesized Berry and Tabor (1977) that, for integrable systems, the levels are uncorrelated, and hence the level spacing distribution follows the one-dimensional Poisson distribution. However for correlated spectrum, the famous Wigner ‘surmise’ Bohr and Mottelson (1969) is a powerful tool to investigate the spacing distribution. In $1984$ Bohigas, Giannoni, and Schmit (BGS) hypothesized Bohigas et al. (1984) that chaotic systems would follow statistical behavior described by random matrix theory appropriate to their symmetry classes. Following BGS, Grobe, Haake, and Sommers (GHS) Grobe et al. (1988) expanded the above spectral contrast between integrability and chaos for Markovian open quantum systems. They found that in the integrable limit, the nearest neighbor spacing distribution of radial distance $s$ for the complex eigenvalues of Liouvillian agree with the two-dimensional (2D) Poisson distribution. Such distribution is the same as Wigner’s conjecture for spacing distribution in the Gaussian orthogonal ensemble (GOE). On the other hand for a fully chaotic limit, the spacing distribution agrees with the Ginibre ensemble  Ginibre (1965); Grobe et al. (1988); Markum et al. (1999); Haake (1991); Hamazaki et al. (2020).
Here we study the eigenvalue properties of the Liouvillian $\mathcal{L}$. Since complex eigenvalues are involved, we consider generalized notions of the level spacing distribution and the consecutive level spacing ratio  Prasad et al. (2022) . By analyzing these quantities as a function of the coupling parameters, we try to sketch the phase diagram which carries information on the non-ergodic to ergodic phase transition in the dissipative ADM.
Figure 6: Scatter plot of the complex spectrum of the Liouvillian $\mathcal{L}$ of the dissipative anisotropic Dicke model for the atom number $N=10$ and $\omega=\omega_{0}=\kappa=1$, for (a) a point in the non-ergodic phase: $g_{1}=1.25,\ g_{2}=0.1$, and (b) a point in the ergodic phase: $g_{1}=1.25,\ g_{2}=1.0$. The bosonic cut-off: $n_{\text{max}}=26$.
The eigenvalues of $\mathcal{L}$ lie in the nonpositive plane, which prevents the divergence of the Eq.(3). The non-real eigenvalues possess complex conjugates, demonstrating the hermiticity of the density matrix. The zero eigenvalue of $\mathcal{L}$ corresponds to a steady state. The real part of the complex eigenvalues corresponds to the loss or decay of the initial density matrix Minganti et al. (2018).
In Fig. 6 we show the scatter plot of the complex spectrum of the Liouvillian $\mathcal{L}$ for two different set of $(g_{1},g_{2})$ pairs: $g_{1}=1.25$ and (a) $g_{2}=0.1$, (b) $g_{2}=1.0$ corresponding to the non-ergodic and ergodic regimes respectively. They clearly show the different natures of the spectrum in the regular and chaotic regimes respectively. For a clearer understanding of the two different phases, one should study the level spacing statistics. Additionally, we would like to discuss briefly how to calculate the level spacing distribution for a non-Hermitian system.
We first carry out an unfolding of the spectrum to remove the system-specific aspects of the level spacing statistics. The complex level spacing distribution $P(s)$ is then produced by using the unfolded spectrum to create a histogram of the Euclidean distance $s$ between nearest-neighbor eigenvalues in the complex plane. For the unfolding process we mainly follow the references Haake (1991); Akemann et al. (2019); Hamazaki et al. (2020).
The Euclidean distance of the nearest neighbor (NN) of each eigenvalue of the Liouvillian is given by, $s_{i}\equiv|E_{i}-E_{i}^{\text{NN}}|$. After rescaling, the analytical expression is given by:
$s_{i}\rightarrow s_{i}^{\prime}=s_{i}\frac{\sqrt{\rho_{\text{av}}(E_{i})}}{\bar{s}}.$ (13)
Here $\rho_{\text{av}}(E_{i})$ is the local average density, which can be approximated in the following way:
$\rho_{\text{av}}(E_{i})=\frac{1}{2\pi\sigma^{2}N}\sum_{i=1}^{N}\exp\left(-\frac{|E-E_{i}|^{2}}{2\sigma^{2}}\right).$ (14)
Following the earlier references Haake (1991); Akemann et al. (2019); Hamazaki et al. (2020), we choose: $\sigma=4.5\times\bar{s}$, where $\bar{s}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$. In Eq 13 $\bar{s}$ is chosen such a way that: $\frac{1}{N}\sum_{i=1}^{N}s_{i}^{\prime}=1$. Finally, we will study the level statistics of the NN distance $s_{i}^{\prime}$.
Figure 7: Level-spacing distribution of the complex spectrum of the Liouvillian $\mathcal{L}$ in (a) the non-ergodic phase with $g_{1}=1.25,\ g_{2}=0.1$ and (b) the ergodic phase with $g_{1}=1.25,\ g_{2}=1.0$. We find remarkable agreement with the 2D Poisson distribution $P_{\text{2D-P}}(s)$ in Eq. and that of the GinUE RMT prediction $P_{\text{GinUE}}(s)$. The atom number $N=10$, bosonic cut-off: $n_{\text{max}}=26$.
The analytical form of $2D$ Poisson distribution, which corresponds to the spacing distribution for independent complex random values is given by:
$P_{2D-P}(s)=\frac{\pi}{2}s\exp\left(-\frac{\pi s^{2}}{4}\right).$ (15)
On the other hand, for the fully chaotic (or ergodic) limit the spacing distribution agrees with the corresponding distribution of the Ginibre ensemble of complex Gaussian non-Hermitian random matrices, given by:
$P_{\text{GinUE}}(s)=\bar{s}\bar{P}_{\text{GinUE}}(\bar{s}s),$ (16)
with
$\bar{P}_{\text{GinUE}}(s)=\sum_{j=1}^{\infty}\frac{2s^{2j+1}\exp(-s^{2})}{\Gamma(1+j,s^{2})}\Pi_{k=1}^{\infty}\frac{\Gamma(1+k,s^{2})}{k!}$ (17)
Figure 8: Scatter plot of the complex level-spacing ratio $z$ introduced in Eq. 18 for the atom number $N=10$, for (a) a point in the non-ergodic phase: $g_{1}=1.25,\ g_{2}=0.1$, which shows the isotropic behavior and (b) a point in the ergodic pgase: $g_{1}=1.25,\ g_{2}=1.0$, which is not isotropic. The bosonic cut-off: $n_{\text{max}}=26$.
[…]
We study the level spacing statistics separately for these two pair points in different phases (see Fig. 7). Fig. 7 displays the striking agreement between the distributions calculated from the spectrum of $\mathcal{L}$ and the $2D$ Poisson distribution in the integrable or regular or the non-ergodic regime and the GinUE forecast in the chaotic or ergodic regime, respectively. In the ergodic phase $P(s)\approx s^{3}$, which indicates the chaotic behavior of the complex eigenvalues as they repel each other and this is consistent with Ginibre distribution in Eq. 16. On the other hand in the non-ergodic phase, we have $P(s)\approx s$ indicating the integrable nature of this regime and consistent with $2D$ Poisson distribution in the complex plane.
Figure 9: Average level spacing ratio, $\langle r\rangle$ of the dissipative anisotropic Dicke model. It shows a transition from the non-ergodic phase (where the system follow the $2D$ Poison statics and $\langle r\rangle\approx 0.69$) to the ergodic phase (where the system behaves as the Ginibre ensemble and $\langle r\rangle\approx 0.74$). Here the atom number $N=8$, the bosonic cut-off $n_{\text{max}}=20$. The other parameters are: $\omega=\omega_{0}=1$, $\kappa=1$.
A.1 Consecutive level spacing ratio
The complex level spacing ratio Sá et al. (2020); Prasad et al. (2022) can be defined as:
$z_{i}=r_{i}e^{\text{i}\theta_{i}}=\frac{E_{i}^{\text{NN}}-E_{i}}{E_{i}^{\text{NNN}}-E_{i}},$ (18)
where the superscripts NN (NNN) stand for nearest (next nearest) neighbor. It captures next-nearest neighbor information that is overlooked in the levelspacing statistics. The advantage of the calculation is we don’t require the unfolding procedure. In Fig. 8 we show the scatter plot of $z$ in the non-ergodic ($g_{1}=1.25,\ g_{2}=0.1$) and the ergodic phases ($g_{1}=1.25,\ g_{2}=1.0$) separately. For the non-ergodic case, the scatter plot is symmetric and isotropic whereas for the ergodic case, it is anisotropic which is another signature for the connection to RMT Sá et al. (2020). In Fig. 9 we show the absolute value of the average level spacing ratio for this model. The figure shows a non-ergodic phase to ergodic phase transition similar to the closed ADM Das et al. (2023a). Here in the non-ergodic phase, $\langle r\rangle\approx 0.69$, (dark color) which is equivalent to the system following the 2D Poisson distribution, whereas in the ergodic phase, $\langle r\rangle\approx 0.74$ (light color), and it corresponds to the Ginibre unitary ensemble.